# TFB: Towards Comprehensive and Fair Benchmarking of Time Series Forecasting Methods

### v1

@@ 1. Introduction
As part of the ongoing digitalization, time series are generated in a variety of domains, such as economic (Sezer et al., 2020; Huang et al., 2022), traffic (Xu et al., 2023; Pedersen et al., 2020; Hu et al., 2018, 2019; Wan et al., 2022; Yang et al., 2023; Hu et al., 2019; Guo et al., 2020; Hu et al., 2017), health (Muspratt and Mammadov, 2023; Miao et al., 2021; Lee et al., 2017; Tran et al., 2019), energy (Alvarez et al., 2010; Guo et al., 2015), and AIOps (Wang et al., 2024; Campos et al., 2022; Wang et al., 2023; Kieu et al., 2022a, b; Zhao et al., 2022). Time Series Forecasting (TSF) is essential in key applications in these domains (Pan et al., 2023; Guo et al., 2014; Yu et al., 2023; Yang et al., 2021). Given historical observations, it is valuable if we can know the future values ahead of time. Correspondingly, TSF has been firmly established as an active research field, witnessing the proposal of numerous methods.
@@
(a) Seasonality
(b) Trend
(c) Shifting
(d) Stationarity
(e) Transition
(f) Non-Seasonality
(g) Non-Trend
(h) Non-Shifting
(i) Non-Stationarity
(j) Non-Transition
@@
Issue 1. Insufficient Coverage of Data Domains. Time series from different domains may exhibit diverse characteristics. Figure 1(a) depicts a time series from the environment domain called AQShunyi (Zhang et al., 2017) that records temperature information at hourly intervals, exhibiting a distinct seasonal pattern. This pattern is reasonable in this scenario because temperatures in nature often cycle around the year. Figure 1(b) shows a time series from FRED-MD (McCracken and Ng, 2016) belongs to economic domain that describes the monthly macroeconomic from 114 regional, national, and international sources with a clear increasing tendency. This may attributed to overall economic stability with minimal fluctuations, reflecting sustained growth in the macroeconomic indicators. Figure 1(c) depicts a series among Electricity (Trindade, 2015) which comes from electricity domain and has a significant change in the data at a certain point in time, which might indicate an abrupt event, etc. However, these simple patterns are only the tip of the iceberg, and time series from different domains may exhibit much more complex patterns that either combine the above characteristics or are entirely different. Therefore, using only limited domains results in limited coverage of time series characteristics, which cannot offer a full picture.
@@
Datasets VAR LR PatchTST NLinear FEDformer Crossformer
NASDAQ 0.462 0.705 0.567 0.522 0.547 0.745
Wind 0.620 0.583 0.652 0.640 0.697 0.590
ILI 1.012 14.244 0.835 0.919 1.020 1.096
Issue 2. Stereotype bias against traditional methods. It is difficult for a single method to exhibit the best performance across all datasets. Methods exhibit varying performance across different datasets. To illustrate the issue, we conduct experiments on three datasets (NASDAQ (Feng et al., 2019), Wind (Li et al., 2022), and ILI (Wu et al., 2021a)) from different domains (stock markets, energy, health) on methods VAR (Toda and Phillips, 1994), PatchTST (Nie et al., 2022), LinearRegression (LR) (Kedem and Fokianos, 2005; Herzen et al., 2022), NLinear (Zeng et al., 2023), FEDformer (Zhou et al., 2022), and Crossformer (Zhang and Yan, 2022). Results are shown in Table 1. Surprisingly, VAR outperforms all recently proposed SOTA methods on NASQAD and is better than FEDfomer and Crossformer on ILI. Furthermore, LR performs better than recently proposed SOTA methods on Wind. However, the experimental studies in their original papers (Zhou et al., 2022; Zhang and Yan, 2022; Nie et al., 2022) do not include VAR and LR in their baselines and rather assume that traditional methods are incapable of obtaining competitive performance. From Table 3, it follows that no existing MTSF benchmark has evaluated statistical methods. Moreover, as the training mechanisms for statistical methods differ from those of deep learning-based methods, it is difficult for existing benchmarks to accommodate statistical methods.
We argue that it is beneficial to eliminate stereotype bias against traditional methods by comparing a wide range of methods.
@@
Table 2. Impact of batch sizes with “drop last.”
Size PatchTST DLinear FEDformer
1 0.4203 0.4874 0.4120
32 0.4138 0.4831 0.4084
64 0.3999 0.4726 0.4022
128 0.3750 0.4539 0.3921
256 0.3561 0.4360 0.3825
512 0.3483 0.4251 0.3736
Issue 3. Lack of consistent and flexible pipelines. The performance of different methods varies with changing experimental settings, e.g., splits among training/validation/testing data, the choice of normalization method, and the selection of hyper-parameter settings. For example, implementations of existing methods often employ a “Drop Last” trick in the testing phase (Zhou et al., 2022; Nie et al., 2022; Wang et al., 2022). To accelerate the testing, it is common to split the data into batches. However, if we discard the last incomplete batch with fewer instances than the batch size, this causes unfair comparisons. For example, in Figure 4, the ETTh2 (Zhou et al., 2021) has a testing series of length 2,880, and we need to predict 336 future time steps using a look-back window of size 512. If we select the batch sizes to be 32, 64, and 128, the number of samples in the last batch are 17, 49, and 113, respectively. Discarding those last-batch testing samples is inappropriate unless all methods use the same strategy. Table 4 shows the testing results of PatchTST (Nie et al., 2022), DLinear (Zeng et al., 2023) and FEDformer (Zhou et al., 2022) on the ETTh2 with different batch sizes and the “Drop Last” trick turned on. We observe that the performance of the methods changes when varying the batch size. However, some existing benchmarks, such as TSlib (Wu et al., 2022), adopt the “Drop Last.” TSlib uses almost the same batch size across all methods in the evaluation phase to achive consistent comparisons. However, when the users select different batch sizes inadvertently, inconsistent comparisons occur. In addition, the pipelines in most benchmarks are inflexible and do not support simultaneous evaluation of statistical learning, machine learning, and deep learning methods. We argue that it is crucial to ensure a consistent and flexible pipeline so that methods are evaluated in the same setting, thereby improving the fairness of the findings.
Robust and extensive benchmarks enable researchers to evaluate proposals for new methods more rigorously across diverse range of datasets, which is crucial for advancing the state-of-the-art (Tan et al., [n.d.]). For example, ImageNet (Deng et al., 2009), that encompasses a substantial dataset has been instrumental in ensuring progress in computer vision. ImageNet has established itself as a standard for assessing methods in image processing due to its support for rigorous evaluation. Table 3 compares existing benchmarks for TSF according to seven properties. No existing benchmark possesses all properties.
@@ 2. RELATED WORK
However, these methods still require manual feature engineering and model design. Taking advantage of the representation learning capabilities offered by deep neural networks (DNNs) on rich data, numerous deep learning methods have been proposed. In many cases, these methods outperform traditional techniques in terms of forecasting accuracy. TCN (Bai et al., 2018) and DeepAR (Salinas et al., 2020) treat time series as sequences of vectors and utilize CNNs or RNNs to capture temporal dependencies. Additionally, Transformer architectures, including Informer (Zhou et al., 2021), FEDformer (Zhou et al., 2022), Autoformer (Wu et al., 2021a), Triformer (Cirstea et al., 2022a), and PatchTST (Nie et al., 2022) can capture more complex temporal dynamics, significantly improving forecasting performance. MLP-based models such as N-HiTS (Challu et al., 2023), N-BEATS (Oreshkin et al., 2019), NLinear (Zeng et al., 2023), and DLinear (Zeng et al., 2023) employ a simple architecture with relatively few parameters and have also demonstrated good forecasting accuracy.
@@ 4.2. Comparison Methods
To investigate the strengths and limitations of different forecasting methods, we include 22 methods that can be categorized as statistical learning, machine learning, and deep learning methods. In terms of statistical learning methods, we include ARIMA (Box and Pierce, 1970), ETS (Hyndman et al., 2008), Kalman Filter (KF) (Harvey, 1990), and VAR (Toda and Phillips, 1994). Among the machine learning methods, we include XGBModel (XGB) (Herzen et al., 2022), LinearRegression (LR) (Kedem and Fokianos, 2005; Herzen et al., 2022), and Random Forest (RF) (Breiman, 2001). Finally, we further split the deep learning methods into RNN-based models (RNN (Kim et al., 2021)), CNN-based models (MICN (Wang et al., 2022), TimesNet (Wu et al., 2022), and TCN (Bai et al., 2018)), MLP-based models (NLinear (Zeng et al., 2023), DLinear (Zeng et al., 2023), TiDE (Das et al., 2023), N-HiTS (Challu et al., 2023), and N-BEATS (Oreshkin et al., 2019)), Transformer-based models (PatchTST (Nie et al., 2022), Crossformer (Zhang and Yan, 2022), and FEDformer (Zhou et al., 2022), Non-stationary Transformer (Stationary) (Liu et al., 2022), Informer (Zhou et al., 2021), and Triformer (Cirstea et al., 2022a)), and Model-Agnostic models (FiLM (Zhou et al., [n.d.])). Figure 6 gives an overview of the diverse technical approaches of the methods covered by the benchmark. This wide selection of comparison methods aims to offer a deeper understanding of the technical nuances and performance variations across methods.
@@ 5.2.2. Multivariate time series forecasting
Figure 9. Comparison of methods across six characteristics (MAE, the outermost edges indicating 0).

### v2

@@ 1. Introduction
As part of the ongoing digitalization, time series are generated in a variety of domains, such as economic (Sezer et al., 2020; Huang et al., 2022), traffic (Xu et al., 2023; Pedersen et al., 2020; Hu et al., 2018; Mo et al., 2022; Wan et al., 2022; Yang et al., 2023; Hu et al., 2019; Guo et al., 2020; Hu et al., 2017), health (Wei et al., 2022; Miao et al., 2021; Lee et al., 2017; Tran et al., 2019), energy (Alvarez et al., 2010; Guo et al., 2015), and AIOps (Wang et al., 2024; Campos et al., 2022; Wang et al., 2023; Qi et al., 2023; Kieu et al., 2022a, b; Zhao et al., 2022). Time Series Forecasting (TSF) is essential in key applications in these domains (Pan et al., 2023; Guo et al., 2014; Yu et al., 2023; Yang et al., 2021). Given historical observations, it is valuable if we can know the future values ahead of time. Correspondingly, TSF has been firmly established as an active research field, witnessing the proposal of numerous methods.
@@
Issue 1. Insufficient Coverage of Data Domains. Time series from different domains may exhibit diverse characteristics. Figure 1a depicts a time series from the environment domain called AQShunyi (Zhang et al., 2017) that records temperature information at hourly intervals, exhibiting a distinct seasonal pattern. This pattern is reasonable in this scenario because temperatures in nature often cycle around the year. Figure 1b shows a time series from FRED-MD (McCracken and Ng, 2016) belongs to economic domain that describes the monthly macroeconomic from 114 regional, national, and international sources with a clear increasing tendency. This may attributed to overall economic stability with minimal fluctuations, reflecting sustained growth in the macroeconomic indicators. Figure 1c depicts a series among Electricity (Trindade, 2015) which comes from electricity domain and has a significant change in the data at a certain point in time, which might indicate an abrupt event, etc. However, these simple patterns are only the tip of the iceberg, and time series from different domains may exhibit much more complex patterns that either combine the above characteristics or are entirely different. Therefore, using only limited domains results in limited coverage of time series characteristics, which cannot offer a full picture.
@@
Datasets
VAR
LR
PatchTST
NLinear
FEDformer
Crossformer
NASDAQ 0.462
0.705
0.567
0.522
0.547
0.745
Wind
0.620
0.583
0.652
0.640
0.697
0.590
ILI
1.012
14.244
0.835 0.919
1.020
1.096
Issue 2. Stereotype bias against traditional methods. It is difficult for a single method to exhibit the best performance across all datasets. Methods exhibit varying performance across different datasets. To illustrate the issue, we conduct experiments on three datasets (NASDAQ (Feng et al., 2019), Wind (Li et al., 2022), and ILI (Wu et al., 2021a)) from different domains (stock markets, energy, health) on methods VAR (Toda and Phillips, 1994), PatchTST (Nie et al., 2022), LinearRegression (LR) (Kedem and Fokianos, 2005; Herzen et al., 2022), NLinear (Zeng et al., 2023), FEDformer (Zhou et al., 2022b), and Crossformer (Zhang and Yan, 2022). Results are shown in Table 1. Surprisingly, VAR outperforms all recently proposed SOTA methods on NASQAD and is better than FEDfomer and Crossformer on ILI. Furthermore, LR performs better than recently proposed SOTA methods on Wind. However, the experimental studies in their original papers (Zhou et al., 2022b; Zhang and Yan, 2022; Nie et al., 2022) do not include VAR and LR in their baselines and rather assume that traditional methods are incapable of obtaining competitive performance. From Table 3, it follows that no existing MTSF benchmark has evaluated statistical methods. Moreover, as the training mechanisms for statistical methods differ from those of deep learning-based methods, it is difficult for existing benchmarks to accommodate statistical methods. We argue that it is beneficial to eliminate stereotype bias against traditional methods by comparing a wide range of methods.
@@
Table 2. Impact of batch sizes with “drop last.” Size PatchTST DLinear FEDformer 1 0.4203 0.4874 0.4120 32 0.4138 0.4831 0.4084 64 0.3999 0.4726 0.4022 128 0.3750 0.4539 0.3921 256 0.3561 0.4360 0.3825 512 0.3483 0.4251 0.3736
Figure 4. “Drop last” illustration.
Issue 3. Lack of consistent and flexible pipelines. The performance of different methods varies with changing experimental settings, e.g., splits among training/validation/testing data, the choice of normalization method, and the selection of hyper-parameter settings. For example, implementations of existing methods often employ a “Drop Last” trick in the testing phase (Zhou et al., 2022b; Nie et al., 2022; Wang et al., 2022). To accelerate the testing, it is common to split the data into batches. However, if we discard the last incomplete batch with fewer instances than the batch size, this causes unfair comparisons. For example, in Figure 4, the ETTh2 (Zhou et al., 2021) has a testing series of length 2,880, and we need to predict 336 future time steps using a look-back window of size 512. If we select the batch sizes to be 32, 64, and 128, the number of samples in the last batch are 17, 49, and 113, respectively. Discarding those last-batch testing samples is inappropriate unless all methods use the same strategy. Table 4 shows the testing results of PatchTST (Nie et al., 2022), DLinear (Zeng et al., 2023) and FEDformer (Zhou et al., 2022b) on the ETTh2 with different batch sizes and the “Drop Last” trick turned on. We observe that the performance of the methods changes when varying the batch size. However, some existing benchmarks, such as TSlib (Wu et al., 2022), adopt the “Drop Last.” TSlib uses almost the same batch size across all methods in the evaluation phase to achive consistent comparisons. However, when the users select different batch sizes inadvertently, inconsistent comparisons occur. In addition, the pipelines in most benchmarks are inflexible and do not support simultaneous evaluation of statistical learning, machine learning, and deep learning methods. We argue that it is crucial to ensure a consistent and flexible pipeline so that methods are evaluated in the same setting, thereby improving the fairness of the findings.
Robust and extensive benchmarks enable researchers to evaluate proposals for new methods more rigorously across diverse range of datasets, which is crucial for advancing the state-of-the-art (Tan et al., 2020). For example, ImageNet (Deng et al., 2009), that encompasses a substantial dataset has been instrumental in ensuring progress in computer vision. ImageNet has established itself as a standard for assessing methods in image processing due to its support for rigorous evaluation. Table 3 compares existing benchmarks for TSF according to seven properties. No existing benchmark possesses all properties.
@@ 2. RELATED WORK
However, these methods still require manual feature engineering and model design. Taking advantage of the representation learning capabilities offered by deep neural networks (DNNs) on rich data, numerous deep learning methods have been proposed. In many cases, these methods outperform traditional techniques in terms of forecasting accuracy. TCN (Bai et al., 2018) and DeepAR (Salinas et al., 2020) treat time series as sequences of vectors and utilize CNNs or RNNs to capture temporal dependencies. Additionally, Transformer architectures, including Informer (Zhou et al., 2021), FEDformer (Zhou et al., 2022b), Autoformer (Wu et al., 2021a), Triformer (Cirstea et al., 2022a), and PatchTST (Nie et al., 2022) can capture more complex temporal dynamics, significantly improving forecasting performance. MLP-based models such as N-HiTS (Challu et al., 2023), N-BEATS (Oreshkin et al., 2019), NLinear (Zeng et al., 2023), and DLinear (Zeng et al., 2023) employ a simple architecture with relatively few parameters and have also demonstrated good forecasting accuracy.
@@ 4.2. Comparison Methods
To investigate the strengths and limitations of different forecasting methods, we include 22 methods that can be categorized as statistical learning, machine learning, and deep learning methods. In terms of statistical learning methods, we include ARIMA (Box and Pierce, 1970), ETS (Hyndman et al., 2008), Kalman Filter (KF) (Harvey, 1990), and VAR (Toda and Phillips, 1994). Among the machine learning methods, we include XGBModel (XGB) (Herzen et al., 2022), LinearRegression (LR) (Kedem and Fokianos, 2005; Herzen et al., 2022), and Random Forest (RF) (Breiman, 2001). Finally, we further split the deep learning methods into RNN-based models (RNN (Kim et al., 2021)), CNN-based models (MICN (Wang et al., 2022), TimesNet (Wu et al., 2022), and TCN (Bai et al., 2018)), MLP-based models (NLinear (Zeng et al., 2023), DLinear (Zeng et al., 2023), TiDE (Das et al., 2023), N-HiTS (Challu et al., 2023), and N-BEATS (Oreshkin et al., 2019)), Transformer-based models (PatchTST (Nie et al., 2022), Crossformer (Zhang and Yan, 2022), and FEDformer (Zhou et al., 2022b), Non-stationary Transformer (Stationary) (Liu et al., 2022), Informer (Zhou et al., 2021), and Triformer (Cirstea et al., 2022a)), and Model-Agnostic models (FiLM (Zhou et al., 2022a)). Figure 6 gives an overview of the diverse technical approaches of the methods covered by the benchmark. This wide selection of comparison methods aims to offer a deeper understanding of the technical nuances and performance variations across methods.
@@ 5.2.2. Multivariate time series forecasting
Figure 9. MAE results of methods across six characteristics.